one of the equal angles in an isosceles triangle is 10 degrees larger than twice the smallest angle in the triangle. What are the degree measures of the triangle?
Answer by mananth(16946)

let the smaller angle be x
the other angle = 2x+10
..
2x+10+2x+10+x=180
5x+20=180
5x=180-20
5x=160
/5
x=160/5
x=32 deg.
the other angles are 2x+10
64+10=74 degrees
...
32,74,74
...
74+74+32=180
...
m.ananth@hotmail.ca